At constant pressure what would be the percentage decrease in density of an ideal gas for a10% increase in temperature

Grade:11

1 Answers

Arun
5 years ago
 
PV=nRT
\rightarrow PV=\frac{Given\ Mass}{Molecular\ Weight}RT
\rightarrow P(M.W)=\frac{Given\ Mass}{V}RT
\rightarrow P(M.W)=\sigma RT                                   where\ \sigma =density
 
Pressure is contant (given), M.W (Molecular Weight) is obviously constant
 
Now Temperature increase is 10%
 
\rightarrow P(M.W)=\sigma (\frac{x}{100})R(\frac{110}{100}T)
\rightarrow (\frac{x}{100})(\frac{110}{100})=1\ (No\ Change)
\rightarrow x=90.9\%
 
Therefore there is a decrease of 9.1% (9 approx)
 
Hope it helps.
 
Regards
Arun (askIITians forum expert)